Find the area under the standard normal curve which lies to the following conditions. Refer to the z-table and to your concept notes for you to be guided.{Hint: To the right means P(x>a), to the left means P(x 1.to the right of z=0.66
2.to the left of z=-1.53
3.to the right of z=-2.34
4.to the right of z=1.30
5.between z=-0.78 and z=0.56

Found 2 solutions by ikleyn, math_tutor2020:
Answer by ikleyn(53937)

.

Do it on your own.

Refer to the z-table and to your concept notes for you to be guided.


        *****************************************

            Why the tutors should do  YOUR  job ?

        *****************************************



Answer by math_tutor2020(3838)


I'll show you how to do problems 1 and 5

Here's the z table I'm using
https://www.ztable.net/
You can use this one, some other similar online resource, or the z table in the back of your book.

===============================================================================
Problem 1

Using that table, you should find that
P(z < 0.66) = 0.74537
which is approximate
Note how we look at the 0.6 row and 0.06 column as shown below

The value at the intersection of the row and column mentioned is what we're after.
In other words, z = 0.66 breaks up into 0.6 + 0.06

From there, we say
P(Z > 0.66) = 1 - P(Z < 0.66)
P(Z > 0.66) = 1 - 0.74537
P(Z > 0.66) = 0.25463

Answer: Approximately 0.25463

Problems 3 and 4 will be handled in a similar fashion (with different numbers of course).

Problem 2 will nearly be identical in steps, but you wont subtract from 1. You just simply need to copy a table value.

For problems 2 and 3, make sure you use the negative z value portion of the table.

===============================================================================

Problem 5

Using that same table, you should find
P(Z < -0.78) = 0.21770
P(Z < 0.56) = 0.71226
both values are approximate

Then we use the formula below
P(a < z < b) = P(z < b) - P(z < a)
P(-0.78 < z < 0.56) = P(z < 0.56) - P(z < -0.78)
P(-0.78 < z < 0.56) = 0.71226 - 0.21770
P(-0.78 < z < 0.56) = 0.49456

Answer: Approximately 0.49456